Alexandra has $15 to buy drinks for her friends at the baseball game. Soda costs $2.75 and bottled water costs $2.00. This relationship can be represented by the inequality . Three of Alexandra's friends asked for water. Which inequality represents the number of sodas she can buy?
Answer by addingup(3677)

Three friends water at $2 each:
2(3)+2.75x = 15
6+2.75x = 15
2.75x = 9
x = 9/2.75 Use your calculator.